\begin{thebibliography}{20}

\bibitem{AKP}B. Alspach, D. L. Kreher and A. Pastine: {\it The Friedlander-Gordon-Miller Conjecture is true,} Austral. J. Combin. {\bf 67}(2017), 11--24.

\bibitem{Anderson}B.A. Anderson, {\it Bailey’s conjecture holds through 87 except possibly for 64,} J. Combin. Math. Combin. Comp. {\bf 12}(1992), 187--195

\bibitem{Anderson3} B.A. Anderson, {\it Sequencings of dicyclic groups}, Ars Combin. {\bf 23}(1987), 131--142. 

\bibitem{AI1}B.A. Anderson and E.C. Ihrig, {\it All groups of odd order have starter-translate 2-sequencings,} Australas. J. Combin. {\bf 6}(1992), 135--146.

\bibitem{AI2} B.A. Anderson and E.C. Ihrig, {\it Every finite solvable group with a unique element of order two, except the quaternion group, has a symmetric sequencing,} J. Combin. Des. {\bf 1}(1993), 3--14.

\bibitem{B} R.A. Bailey, {\it Quasi-complete latin squares: construction and randomization}, J. R. Stat. Soc. Ser. B Stat. Methodol. {\bf 46}(1984), 323--334.

\bibitem{JG} Beals, R. and Gallian, J.A. and Headley, P. and Jungreis, D, {\it Harmonious groups}, Journal of Combinatorial Theory, Series A {\bf 56}(1991), no. 2, 223--238.

\bibitem{Bray} J.N. Bray, Q. Cai, P.J. Cameron, P. Spiga, H. Zhang, {\it The Hall–Paige conjecture, and synchronization for affine and diagonal groups},  J. Algebra {\bf 545}(2020), 27--42.

\bibitem{DJ} L. De Wolf and M. Javaheri, {\it Harmonious sequences in groups with a unique involution}, Combin. Theory {\bf 5}(2025), no. 3, Article 65556.

\bibitem{Evans} A. Evans, {\it The admissibility of sporadic simple groups}, J. Algebra {\bf 321}(2009), no. 1, 105--116.

\bibitem{G} B. Gordon, {\it Sequences in groups with distinct partial products,} Pacific J. Math. {\bf 11}(1961), no. 4, 1309--1313.

\bibitem{HP} M. Hall and L. Paige, {\it Complete mappings of finite groups}, Pacific J. Math. {\bf 5}(1955), 541--549. 

\bibitem{Isbell} J. Isbell, {\it Sequencing certain dihedral groups,} Discrete Math. {\bf 85}(1990), 323--328. 

\bibitem{Jav} M. Javaheri, {\it Doubly sequenceable groups}, J. Combin. Des. {\bf 32}(2024), no. 7, 371--387.

\bibitem{K2} A.D. Keedwell, {\it On the R-sequenceability and $R_h$-sequenceability of groups}, Ann. Discrete Math. {\bf 18}(1983), 535--548.

\bibitem{Keedwell}A.D. Keedwell, {\it Sequenceable Groups, Generalized Complete Mappings, Neofields and Block Designs, Proc. of Tenth Austral. Conf. on Comb. Math., Lecture Notes in Math. \#1036 (1983), 49--71, Springer-Verlag.}

\bibitem{Li}P. Li, {\it Sequencing the dihedral groups $D_{4k}$,} Discrete Math. {\bf 175} (1997) 271--276.

\bibitem{MP} A. M\"{u}esser and A. Pokrovskiy, {\it A random {H}all-{P}aige conjecture}, Invent. Math. {\bf 240}(2025), 779--867.

\bibitem{Ollis3} M.A. Ollis, {\it A note on terraces for abelian groups}, Australas. J. Combin. {\bf 52}(2012), 229--234.

\bibitem{O1} M.A. Ollis, {\it On terraces for abelian groups}, Discrete Math. {\bf 305}(2005), 250--263.

\bibitem{Ollis} M.A. Ollis, {\it Sequenceable groups and related topics}, Electron. J. Combin., DS10-May.

\bibitem{OW} M.A. Ollis and D.T. Willmott, {\it Constructions for terraces and R-sequencings, including a proof that Bailey’s Conjecture holds for abelian groups}, J. Combin. Des. {\bf 23}(2015), 1--17.

\bibitem{OW2} M.A. Ollis and D.T. Willmott, {\it On twizzler, zigzag and graceful terraces}, Australas. J. Combin. {\bf 51}(2011), 243--257.

\bibitem{Paige} L.J. Paige, {\it Complete mappings of finite groups}, Pacific J. Math. {\bf 1}(1951), 111--116.

\bibitem{Wang} C.-D. Wang, {\it On the harmoniousness of dicyclic groups}, Discrete Math. {\bf 120}(1993), no. 1, 221--225.

\bibitem{WL} C.-D. Wang and P.A. Leonard, {\it More on sequences in groups,} Australas. J. Combin. {\bf 21}(2000), 187--196.

\bibitem{Wilcox} S. Wilcox, {\it Reduction of the Hall-Paige conjecture to sporadic simple groups}, J. Algebra {\bf 321}(2009), no. 5, 1407--1428.

\end{thebibliography}

\documentclass[10pt]{article}
\usepackage{amssymb,float}

\usepackage[margin=1in]{geometry}
\usepackage{amsthm}
\usepackage{amsmath,mathabx,pifont}
\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{cor}[theorem]{Corollary}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{example}[theorem]{Example}

\theoremstyle{remark}

\usepackage{multirow}
\newcommand\mcl[1]{\multicolumn{2}{|l|}{#1}}

\numberwithin{equation}{section}
\title{Symmetric sequencings and other combinatorial properties \\ of large groups}
\author{ Mohammad Javaheri  \\
515 Loudon Road \\
Siena University, School of Science\\
Loudonville, NY 12211, USA\\
\small{mjavaheri@siena.edu}  
}

\begin{document}

\maketitle

\begin{abstract}
We prove that Anderson's conjecture on symmetric sequencings and Bailey's conjecture on 2-sequencings hold for sufficiently large groups. In addition, we discuss extensions of partial harmonious sequences and partial R-sequencings. Several further results on double sequencings are presented, both in the context of abelian groups and for sufficiently large non-abelian groups.
\end{abstract}

\section{Introduction}
In 1961, motivated by the study of Latin squares, Gordon defined the notion of a sequenceable group \cite{G}. A finite group $G$ of order $n > 1$ is {\it sequenceable} if there exists an ordering $b_1, b_2, \dots, b_{n}$ of its elements with $b_1=1_G$, the identity element of $G$, such that the partial products $a_i = b_1 \cdots b_i, 1\leq i  \leq n$, are all distinct. Such sequencings provide a constructive method to generate row-complete Latin squares with applications in combinatorial design theory.

Gordon \cite{G} also proved that a finite abelian group is sequenceable if and only if it is binary (having a unique element of order 2). He also established that the dihedral groups of orders 6 and 8, are not sequenceable. Subsequently, Isbell and Li \cite{Isbell,Li} proved all dihedral groups of order at least $10$ are sequenceable, and Anderson \cite{AI2} proved that every solvable binary group, except the quaternion group, is sequenceable. 

These results, together with extensive computational evidence, underpin Keedwell’s conjecture \cite{Keedwell}: Every non-abelian group of order at least 10 is sequenceable. The strongest confirmation of Keedwell's conjecture was provided recently by M\"{u}esser and Pokrovskiy \cite{MP} who prove that all sufficiently large groups are sequenceable. We use their results to establish several other combinatorial properties of sufficiently large groups. 

Let ${\bf g}: g_1,\ldots, g_{m}$ be a sequence in a group $G$. For $1\leq i \leq m$, let 
$$\bar g_i=g_{i-1}^{-1}g_i \quad \hat g_i=g_{i-1}g_i,$$ 
where $g_{0}=g_m$. We define three derived sequences and several related combinatorial properties related to sequenceability. Given a sequence $g_1,g_2,\ldots, g_m$ in $G$, let
\begin{align*}
{\bf \bar g}&:  g_1, \bar{g}_2,\ldots, \bar{g}_{m},\\
{\bf \check g}&:  \bar{g}_1,\bar{g}_2,\ldots, \bar{g}_{m},\\
{\bf \hat g}&:  \hat{g}_1, \hat{g}_2,\ldots, \hat{g}_{m}.
\end{align*}

\begin{definition}
For a subset $A$ of a finite group $G$, let ${\cal P}_k(A)$ denote the set of sequences in $A$ that contain every element of $A$ exactly $k$ times. 
 \begin{itemize}
 \item[i)] $A$ is {\it sequenceable} if $\exists {\bf g} \in {\cal P}_1(A)$ such that $g_1=1_G$ and ${\bf \bar g} \in {\cal P}_1(A)$, in which case ${\bf \bar g}$ is called a {\it sequencing} of $G$. It is {\it doubly sequenceable} if $\exists {\bf g} \in {\cal P}_2(A)$ such that $g_1=1_G$ and ${\bf \bar g} \in {\cal P}_2(A)$.
  
 \item[ii)] $A$ is {\it R-sequenceable} if $\exists {\bf g} \in {\cal P}_1(A\setminus \{1_G\})$ such that ${\bf \check g} \in {\cal P}_1(A\setminus \{1_G\})$, in which case ${\bf \check g}$ is called an {\it R-sequencing} of $A$. It is {\it doubly R-sequenceable} if $\exists {\bf g} \in {\cal P}_2(A\setminus \{1_G\})$ such that ${\bf \check g} \in {\cal P}_2(A\setminus \{1_G\})$
 
 \item[iii)] $A$ is {\it symmetric sequenceable} if $\exists {\bf g} \in {\cal P}_1(A)$ with $g_1=1_G$ such that ${\bf \bar g} \in {\cal P}_1(A)$ and $\bar{g}_i\bar{g}_{m+2-i}=1_G$ for all $1\leq i \leq m$, where $|A|=m$. 
 
 \item[iv)] $A$ is {\it $2$-sequenceable} if there exists a sequence $g_1,\ldots, g_m$ of elements of $A$ with $g_1=1_G$ such that ${\bf \bar g} \in {\cal P}_1(A)$ and for each element $g\in A$:
 \begin{align*}
 g^2 \neq 1_G & \Rightarrow \#\{1\leq i \leq m: \bar{g}_i=g~\mbox{\it or}~\bar{g}_i=g^{-1}\}=2,\\
  g^2 = 1_G & \Rightarrow \#\{1\leq i \leq m: \bar{g}_i=g \}=1.
 \end{align*}

 \item[v)] $A$ is {\it harmonious} if $\exists {\bf g} \in {\cal P}_1(A)$ such that ${\bf \hat g} \in {\cal P}_1(A)$, in which case ${\bf g}$ is called a {\it harmonious sequence} in $A$. It is {\it doubly harmonious} if $\exists {\bf g} \in {\cal P}_2(A)$ such that ${\bf \hat g} \in {\cal P}_2(A)$. 
 
 \item [v)] $A$ is {\it R-harmonious} if $A\setminus \{1_G\}$ is harmonious, and it is {\it doubly R-harmonious} if $A\setminus \{1_G\}$ is doubly harmonious. 
 \end{itemize}
\end{definition}

 Harmonious sequences and R-sequencings of a group $G$ naturally give rise to {\it complete mappings} (a bijection $\phi: G \rightarrow G$ such that $x\mapsto x \phi(x)$ is also a bijection). Hall and Paige \cite{HP} proved that a necessary condition for a group to admit a complete mapping is that its Sylow $2$-subgroups are either trivial or non-cyclic (now known as the Hall-Paige condition). They conjectured that this condition is also sufficient. The Hall-Paige conjecture was eventually confirmed in 2009 by Bray, Evans, and Wilcox \cite{Bray,Evans,Wilcox}. 

\begin{center}
    \begin{table}
  %  \footnotesize
    \renewcommand\arraystretch{1.2}
      \centering
\begin{tabular}{|*{4}{l|l|c|c|c|c|}}
  \hline
\mcl{\textbf{}}      
    &   \textbf{S}    &   \textbf{RS}     & \textbf{H} & \textbf{RH} & \textbf{SS} & \textbf{2S} & \textbf{D}
        \\ \hline \hline
\multirow{3}{15mm}{Abelian order $\geq 4$}

& binary $\neq \mathbb{Z}_2$
& \checkmark & \ding{53}    & \ding{53} & \ding{53} & \checkmark &\checkmark & \checkmark
        \\
  \cline{2-9} 
    &   $(\mathbb{Z}_2)^n,~n\geq 2$ & \ding{53} &  \checkmark   & \ding{53}  &\checkmark & \ding{53} & \ding{53}  &\checkmark
            \\
       \cline{2-9} 
    &  all other $\neq \mathbb{Z}_3$ &   \ding{53}  &   \checkmark    & \checkmark    & \checkmark   & \ding{53}    &\checkmark  &\checkmark 
                       \\                  \hline \hline

  \multirow{4}{20mm}{Binary\\ order $\geq 10$}  
  
  &   $Q_{8n}$, $n \geq 2$  &\checkmark &  \checkmark  & \checkmark &     & \checkmark  & \checkmark  &C     \\ 
\cline{2-9}
    &  $\mathbb{Z} \Lambda$    &   \checkmark    &    \ding{53}       & \ding{53}   &\ding{53}   & \checkmark   & \checkmark    &C
                       \\     
                       \cline{2-9}
    &  solvable $Q \Lambda$ $\neq Q_8$    &   \checkmark   &      &   & & \checkmark & \checkmark &C        \\     
    \cline{2-9}

    & non-solvable $Q \Lambda$
        &   C    &                &   &   & C  & C &C
                       \\     \hline \hline

 \multirow{1}{20mm}{Dihedral \\ order $\geq 10$}
     
    &   order $4n\geq 12$    & \checkmark  &  \checkmark   &    &   & \ding{53}   & \checkmark   & C
            \\
          \cline{2-9} 
    &  order $4n+2 \geq 10$   &  \checkmark &   \ding{53}  & \ding{53}  &\ding{53}  & \ding{53}  &\checkmark  &C
                       \\     \hline \hline

\mcl{Non-abelian odd-order}  
        & C     &  &  \checkmark   &  &\ding{53}  &\checkmark  &C
               \\      \hline \hline
\multirow{3}{20mm}{Non-abelian large order}
    &   HP and binary   & \checkmark  &  \checkmark   & \checkmark  &\checkmark   &\checkmark & \checkmark  & \checkmark
            \\   
         \cline{2-9} 
    &  binary not HP   &   \checkmark   &   \ding{53}             & \ding{53}   & \ding{53}   & \checkmark & \checkmark  & \checkmark
 \\
         \cline{2-9} 
    & HP not binary  &   \checkmark &   \checkmark           & \checkmark & \checkmark & \ding{53} & \checkmark & \checkmark
\\
  \cline{2-9} 
    &  not HP not binary   &   \checkmark   &   \ding{53}    & \ding{53}  & \ding{53}  & \ding{53}  & \checkmark  & \checkmark
                       \\     
     \hline
\end{tabular}
\label{tab1}
\caption{ {\bf S}: Sequenceable, {\bf RS}: R-sequenceable, {\bf H}: harmonious, {\bf RH}: R-harmonious, {\bf SS}: symmetric sequenceable. {\bf 2S}: 2-sequenceable, {\bf D}: doubly sequenceable, doubly R-sequenceable, doubly harmonious, and doubly R-harmonious, {\bf $\mathbb{Z}\Lambda$}: binary groups whose Sylow 2-subgroups are cyclic, {\bf $Q\Lambda$}: binary groups whose Sylow 2-subgroups are non-cyclic, $Q_{8n}=\langle a,b: a^{4n}=e,b^2=a^{2n}, ab=ba^{-1} \rangle$ denotes the dicyclic group of order $8n$,  {\bf HP}: the group satisfies the Hall–Paige condition, {\bf C}: a conjecture exists that the answer is affirmative for the corresponding class/property. Empty cells have no explicit conjecture recorded in the literature, but the property is widely expected to hold.}
    \end{table}
 \end{center}

Table \ref{tab1} and the following discussion summarize many of the known results on sequenceability and several related sequencing properties, incorporating the contributions of the present article. For a comprehensive and regularly updated survey of sequenceable groups, 2-sequencings (terraces), R-sequencings, harmonious sequences, and related structures, the reader is referred to Ollis~\cite{Ollis}.
 
\begin{itemize}
\item {\bf Sequenceability.} Sequenceable groups include all solvable binary groups (including abelian groups) except the quaternion group $Q_8$ \cite{AI2,G}, dihedral groups of order $2n\geq 10$ \cite{Isbell, Li}, and all sufficiently large non-abelian groups \cite{MP}.

\item {\bf Symmetric sequenceability.} Symmetric sequenceable groups include all solvable binary groups except $Q_8$ \cite{AI2}. Anderson stated the following conjecture in 1993:

\medskip

\noindent {\bf Conjecture 1}. (Anderson \cite{Anderson}) {\it All binary groups except $Q_8$ are symmetric sequenceable. }

\medskip

We prove that if $G$ is a sufficiently large binary group, then $G$ is symmetric sequenceable (Cor.\ \ref{ss}).

\item {\bf R-sequenceability.} R-sequenceable groups include all non-binary abelian groups \cite{AKP}, dihedral groups of order $4n\geq 4$ \cite{K2}, dicyclic groups of order $8n\geq 16$ \cite{WL}, and all sufficiently large groups that satisfy the Hall-Paige condition \cite{MP}.

\item {\bf 2-sequenceability.} 2-sequenceable groups include all abelian groups except the elementary 2-groups of order at least 4 \cite{Ollis3,O1,OW,OW2}, and all non-abelian odd-order groups \cite{AI1}. Bailey stated the following conjecture in 1984:

\medskip

\noindent {\bf Conjecture 2.} (Bailey \cite{B}) {\it All finite groups, except the elementary $2$-groups of order at least $4$, are $2$-sequenceable.}

\medskip

We prove that if $G$ is a sufficiently large group that is not an elementary 2-group, then $G$ is $2$-sequenceable (Cor.\ \ref{2s}).  

\item {\bf Harmoniousness.} Harmonious groups include all odd-order groups, all abelian groups except the elementary 2-groups and binary groups, dihedral groups of order $8n\geq 8$ \cite{JG}, dicyclic groups of order $8n\geq 16$ \cite{Wang}, and all sufficiently large groups that satisfy the Hall-Paige condition \cite{MP}.

If $G$ is a non-binary abelian group and $G\neq \mathbb{Z}_3$, then $G$ is R-harmonious \cite{JG} (Beals et al.\ used the term $\#$-harmonious). If $G$ is a binary abelian group with a unique element $\imath_G$ of order 2, then $G\backslash \{\imath_G\}$ is harmonious \cite{DJ}. See Corollary \ref{bun} for similar results in sufficiently large groups. 

\item {\bf Double sequences.} All abelian groups are doubly sequenceable \cite{Jav} and doubly harmonious \cite{DJ}. If $G$ is an abelian group of order at least 4, then $G$ is doubly R-harmonious \cite{DJ}. We prove that every abelian group except $\mathbb{Z}_2$ is doubly R-sequenceable (Theorem \ref{drs}). This completes the investigation in the abelian case of the conjectures stated in \cite{DJ,Jav}:

\medskip

\noindent {\bf Conjecture 3}. {\it All groups of order at least 4 are doubly sequenceable, doubly R-sequenceable, doubly harmonious, and doubly R-harmonious. }

\medskip

In addition, we verify this conjecture for sufficiently large groups in Theorem \ref{bun2}.

\end{itemize}

In Section 2, we prove that all abelian groups except $\mathbb{Z}_2$ are doubly R-sequenceable (Theorem \ref{drs}). In Section 3, we verify Anderson's and Bailey's conjectures for sufficiently large groups (Cor.\ \ref{2} and \ref{ss}). We also discuss supersequenceable groups and prove that all sufficiently large groups are supersequenceable (Cor.\ \ref{superseq}). In Section \ref{4}, we discuss extensions of partial harmonious sequences and partial R-sequencings and prove several results on sufficiently large groups that complement the existing results in abelian groups. 

\section{DR-sequenceable groups}\label{2}

In this section, we prove that all abelian groups except $\mathbb{Z}_2$ are doubly R-sequenceable (DR-sequenceable). Every abelian group that is not binary is R-sequenceable, hence DR-sequenceable. Thus, suppose that $G=\mathbb{Z}_{2^k} \times H$, where $H$ is an odd-order abelian group and $k\geq 1$. We first consider the case $k=1$. 

A {\it symmetric harmonious} sequence in a group $H$ of order $n$ is a harmonious sequence $h_0,\ldots, h_{n-1}$ of $H$ with $h_0=1_H$ such that 
$$h_ih_{n-i}=1_H~\mbox{for all}~1\leq i \leq n-1.$$

If $H$ is an odd-order group, then $\mathbb{Z}_2 \times H$ is symmetric sequenceable \cite{AI2} (since it is solvable and binary) and doubly harmonious \cite{DJ}, but not harmonious, R-sequenceable, or R-harmonious (since it does not satisfy the Hall-Paige condition). We next prove that such groups are doubly R-sequenceable. To illustrate the proof idea, consider the example of $G=\mathbb{Z}_2 \times \mathbb{Z}_5$. To obtain a DR-sequencing of $G$, we start with a symmetric harmonious sequence ${\bf h}: 0,1,2,3,4$ in $\mathbb{Z}_5$. Then double the sequence and shuffle in the additive inverses to obtain:
$$\underline{1},4,\underline{2},3,\underline{3},2,\underline{4},1,0,\underline{1},4,\underline{2},3,\underline{3},2,\underline{4},1,0.$$
Next, we pair each number with $0$ or $1$ according to the pattern below, where the element $(a,b) \in \mathbb{Z}_2 \times \mathbb{Z}_5$ is denoted by $\begin{matrix}a \\ b \end{matrix}$. 
$$\overbrace{\begin{matrix}1 \\ 1 \end{matrix},\begin{matrix}1 \\ 4 \end{matrix},\begin{matrix}0 \\ 2 \end{matrix},\begin{matrix}0 \\ 3 \end{matrix},\begin{matrix}1 \\ 3 \end{matrix},\begin{matrix}1 \\ 2 \end{matrix},\begin{matrix}0 \\ 4 \end{matrix},\begin{matrix}0 \\ 1 \end{matrix}}^\text{$1100$},\begin{matrix}1 \\ 0 \end{matrix},\overbrace{\begin{matrix}0 \\ 1 \end{matrix},\begin{matrix}1 \\ 4 \end{matrix},\begin{matrix}1 \\ 2 \end{matrix},\begin{matrix}0 \\ 3 \end{matrix}}^\text{$0110$},\overbrace{\begin{matrix}1 \\ 3 \end{matrix},\begin{matrix}0 \\ 2 \end{matrix},\begin{matrix}0 \\ 4 \end{matrix},\begin{matrix}1 \\ 1 \end{matrix}}^\text{$1001$},\begin{matrix}1 \\ 0 \end{matrix},$$
and the consecutive differences are
$$\begin{matrix}0 \\ 1 \end{matrix},\begin{matrix}0 \\ 3 \end{matrix},\begin{matrix}1 \\ 3 \end{matrix},\begin{matrix}0 \\ 1 \end{matrix},\begin{matrix}1 \\ 0 \end{matrix},\begin{matrix}0 \\ 4 \end{matrix},\begin{matrix}1 \\ 2 \end{matrix},\begin{matrix}0 \\ 2 \end{matrix},\begin{matrix}1 \\ 4 \end{matrix},\begin{matrix}1 \\ 1 \end{matrix},\begin{matrix}1 \\ 3 \end{matrix},\begin{matrix}0 \\ 3 \end{matrix},\begin{matrix}1 \\ 1 \end{matrix},\begin{matrix}1 \\ 0 \end{matrix},\begin{matrix}1 \\ 4 \end{matrix},\begin{matrix}0 \\ 2 \end{matrix},\begin{matrix}1 \\ 2 \end{matrix},\begin{matrix}0 \\ 4 \end{matrix},$$
which is a DR-sequencing of $\mathbb{Z}_2 \times \mathbb{Z}_5$. 

\begin{theorem}\label{DR-2}
If $H$ is a nontrivial odd group, then $G=\mathbb{Z}_2 \times H$ is DR-sequenceable. 
\end{theorem}

\begin{proof}
Let ${\bf h}:h_0,h_1,\ldots, h_{n-1}$ be a symmetric harmonious sequence in $H$, where $n=|H|$. Define the sequence $S_1$ by 
$$S_1:~(1,h_1),(1,h_1^{-1}),(0,h_2),(0,h_2^{-1}),\ldots, (0,h_{n-1}),(0,h_{n-1}^{-1}),$$
where the $i$th pair in $S_1$ is given by $(1,h_i),(1,h_i^{-1})$ if $i$ is odd, and it is given by $(0,h_i),(0,h_i^{-1})$ if $i$ is even. Define the sequence $S_2$ by 
$$S_2:~(0,h_1),(1,h_1^{-1}),(1,h_2),(0,h_2^{-1}),\ldots, (0,h_{n-1}),(1,h_{n-1}^{-1}),$$
where the $i$th pair in $S_2$ is given by $(0,h_i), (1,h_i^{-1})$ if $i<n/2$ is odd and $i>n/2$ is even, and it is given by $(1,h_i),(0,h_i^{-1})$ if $i<n/2$ is even or $i>n/2$ is odd. We claim that the sequence 
$${\bf g}:~S_1, (1,h_0), S_2, (1,h_0)$$
is a DR-sequencing in $G$. 

To be more precise, given $0\leq i \leq 4n-3$, let $j=\lceil i/2 \rceil +1 \pmod n$. Define
$$g_i=\begin{cases}
(1,h_{j}) & \mbox{if $0\leq i\leq 2n-2$ and $i=0 \pmod 4$};\\
(1,h_{j}^{-1}) & \mbox{if $0\leq i\leq 2n-2$ and $i=1 \pmod 4$};\\
(0,h_{j}) & \mbox{if $0\leq i\leq 2n-2$ and $i=2 \pmod 4$};\\
(0,h_{j}^{-1}) & \mbox{if $0\leq i\leq 2n-2$ and $i=3 \pmod 4$};\\
(0,h_{j}) & \mbox{if $2n-2< i\leq 3n-3$ and $i=1 \pmod 4$};\\
(1,h_{j}^{-1}) & \mbox{if $2n-2< i\leq 3n-3$ and $i=2 \pmod 4$};\\
(1,h_{j}) & \mbox{if $2n-2< i\leq 3n-3$ and $i=3 \pmod 4$};\\
(0,h_{j}^{-1}) & \mbox{if $2n-2< i\leq 3n-3$ and $i=0 \pmod 4$};\\
(1,h_{j}) & \mbox{if $3n-3< i\leq 4n-3$ and $i=1 \pmod 4$};\\
(0,h_{j}^{-1}) & \mbox{if $3n-3< i\leq 4n-3$ and $i=2 \pmod 4$};\\
(0,h_{j}) & \mbox{if $3n-3< i\leq 4n-3$ and $i=3 \pmod 4$};\\
(1,h_{j}^{-1}) & \mbox{if $3n-3< i\leq 4n-3$ and $i=0 \pmod 4$};\\
\end{cases}
$$
It follows that:
$$\bar g_i=\begin{cases}
(0,h_0h_1) & \mbox{if $i=0$};\\
(1,h_{j-1}h_{j}) & \mbox{if $0< i\leq 2n-2$ and $i=0 \pmod 4$};\\
(0,h_{j}^{-2}) & \mbox{if $0\leq i\leq 2n-2$ and $i=1 \pmod 4$};\\
(1,h_{j-1}h_{j}) & \mbox{if $0\leq i\leq 2n-2$ and $i=2 \pmod 4$};\\
(0,h_{j}^{-2}) & \mbox{if $0\leq i\leq 2n-2$ and $i=3 \pmod 4$};\\
(0,h_{j-1}h_{j}) & \mbox{if $2n-2< i\leq 3n-3$ and $i=1 \pmod 4$};\\
(1,h_{j}^{-2}) & \mbox{if $2n-2< i\leq 3n-3$ and $i=2 \pmod 4$};\\
(0,h_{j-1}h_{j}) & \mbox{if $2n-2< i\leq 3n-3$ and $i=3 \pmod 4$};\\
(1,h_{j}^{-2}) & \mbox{if $2n-2< i\leq 3n-3$ and $i=0 \pmod 4$};\\
(0,h_{j-1}h_{j}) & \mbox{if $3n-3< i\leq 4n-3$ and $i=1 \pmod 4$};\\
(1,h_{j}^{-2}) & \mbox{if $3n-3< i\leq 4n-3$ and $i=2 \pmod 4$};\\
(0,h_{j-1}h_{j}) & \mbox{if $3n-3< i\leq 4n-3$ and $i=3 \pmod 4$};\\
(1,h_{j}^{-2}) & \mbox{if $3n-3< i\leq 4n-3$ and $i=0 \pmod 4$};\\
\end{cases}
$$
It is straightforward to check that $\bar g_0,\bar g_1,\ldots, \bar g_{4n-3}$ is a DR-sequencing of $G$. \end{proof}

For the case $G=\mathbb{Z}_{2^k} \times H$ and $k \geq 2$, we will need a sequencing of $G$ where the second term is of a specific form. The following theorem proves a more general statement. 
 
We can first demonstrate the idea of the proof by an example. To construct a symmetric sequencing of $\mathbb{Z}_4 \times \mathbb{Z}_3$, we start with the sequence ${\bf k}:0,1,3,2$ which gives the symmetric sequencing ${\bf \bar k}: 0, 1, 2, 3$ in $\mathbb{Z}_4$. Then repeat the sequence ${\bf k}$ three times in alternating forms of regular and reverse order:
$$\underbrace{0,1,3,2}_\text{regular},\underbrace{2,3,1,0}_\text{reverse},\underbrace{0,1,3,2}_\text{regular}.$$
Next, we pair the terms of a symmetric harmonious sequence ${\bf h}:0,1,2$ in $\mathbb{Z}_3$ with the half-blocks as follows:
$${\bf g}:\overbrace{\begin{matrix} 0 \\ 0 \end{matrix}, \begin{matrix} 0 \\ 1 \end{matrix}}^\text{00},\overbrace{\begin{matrix} 1 \\ 3 \end{matrix}, \begin{matrix} 2 \\ 2 \end{matrix},\begin{matrix} 1 \\ 2 \end{matrix},\begin{matrix} 2 \\ 3 \end{matrix}}^\text{12}, \overbrace{\begin{matrix} 2 \\ 1 \end{matrix},\begin{matrix} 1 \\ 0 \end{matrix}, \begin{matrix} 2 \\ 0 \end{matrix}, \begin{matrix} 1 \\ 1 \end{matrix}}^\text{21},\overbrace{\begin{matrix} 0 \\ 3 \end{matrix},\begin{matrix} 0 \\ 2 \end{matrix}}^\text{00},
$$
and the consecutive differences are:
$${\bf \bar g}:\begin{matrix} 0 \\ 0 \end{matrix}, \begin{matrix} 0 \\ 1 \end{matrix},\begin{matrix} 1 \\ 2 \end{matrix}, \begin{matrix} 1 \\ 3 \end{matrix},\begin{matrix} 2 \\ 0 \end{matrix},\begin{matrix} 1 \\ 1 \end{matrix}, \begin{matrix} 0 \\ 2 \end{matrix},\begin{matrix} 2 \\ 3 \end{matrix}, \begin{matrix} 1 \\ 0 \end{matrix}, \begin{matrix} 2 \\ 1 \end{matrix},\begin{matrix} 2 \\ 2 \end{matrix},\begin{matrix} 0 \\ 3 \end{matrix},
$$
which is a symmetric sequencing of $\mathbb{Z}_4 \times \mathbb{Z}_3$. 

\begin{theorem}\label{newseqgen}
Let $G=K \times H$, where $K$ is a symmetric sequenceable group and $H$ is an odd group. Then $G$ is symmetric sequenceable. 
\end{theorem}

\begin{proof}
Since $K$ is symmetric sequenceable, there exists ${\bf k}:1_K=k_0,\ldots, k_{n-1}$, a permutation of elements of $K$, such that ${\bf \bar k}: 1_K=k_0, \bar k_1,\ldots, \bar k_{n-1}$ is also a permutation of elements of $K$ and 
$$\bar k_i  \bar k_{n-i}=1_K~\mbox{for all}~1\leq i \leq n-1.$$
Let $\alpha_0,\alpha_1,\ldots, \alpha_{m-1}$ be a symmetric harmonious sequence in $H$ (with $\alpha_m=\alpha_0=1_H$). Given $0\leq i \leq mn-1$, write $i=qn+r$, where $0\leq q \leq m-1$ and $0\leq r \leq n-1$.
We define a sequence ${\bf g}:~g_0,\ldots, g_{mn-1}$ in $G$ by letting  $s=(-1)^r$ and
\begin{equation}\label{g11}
g_i=\begin{cases}
\left ( k_r , \alpha_q^s \right )& \mbox{if $q$ is even and $0\leq r<n/2$},\\
\left ( k_r , \alpha_{q+1}^s\right )& \mbox{if $q$ is even and $n/2 \leq r \leq n-1$},\\
\left ( k_{n-1-r}, \alpha_q^s \right ) & \mbox{if $q$ is odd and $0\leq r <n/2$},\\
\left ( k_{n-1-r}, \alpha_{q+1}^s \right ) & \mbox{if $q$ is odd and $n/2 \leq r \leq n-1$}.
\end{cases}
\end{equation}
It follows that
$$\bar g_i =\begin{cases}
\left (1_K,\alpha_q^{2} \right ) & \mbox{if $r=0$},\\
\left (\bar k_r,\alpha_q^{2s} \right ) & \mbox{if $0< r<n/2$},\\
\left (\bar k_r,\alpha_q \alpha_{q+1} \right ) & \mbox{if $r=n/2$},\\
\left ( \bar k_r , \alpha_{q+1}^{2s}\right )& \mbox{if $n/2 < r \leq n-1$}.\\
\end{cases}
$$
It is straightforward to check that ${\bf \bar g}$ is a sequencing of $G$. To see it is a symmetric sequencing, given $1\leq i \leq mn-1$, write $i=qn+r$, where $0\leq q \leq m-1$ and $0\leq r \leq n-1$. Consider the case where $0<r<n/2$. Then $mn-i=mn-qn-r=(m-q-1)n+n-r$, and so
$${\bar g_i}{\bar g_{mn-i}}=(\bar k_r, \alpha_q^{2s})\cdot (\bar k_{n-r},\alpha_{m-q}^{2s})=(1_K,1_H)=1_G.$$
If $r=n/2$, then 
$${\bar g_i}{\bar g_{mn-i}}=(\bar k_r, \alpha_q \alpha_{q+1})\cdot (\bar k_{n-r}, \alpha_{m-q-1}\alpha_{m-q})=1_G.$$
The proof in other cases is similar, and it follows that $G$ is symmetric sequenceable. 
\end{proof}

We are now ready to prove the main theorem of this section.   

\begin{theorem}\label{drs}
Every abelian group except $\mathbb{Z}_2$ is DR-sequenceable. 
\end{theorem}

\begin{proof}
If $G$ is not binary, then $G$ is R-sequenceable \cite{AKP}, hence DR-sequenceable. Thus, suppose that $G=\mathbb{Z}_{2^k} \times H$, where $H$ is an odd group. The case where $k=1$ and $H$ is nontrivial follows from Theorem \ref{DR-2}. Thus, suppose that $k\geq 2$. Let $n=2^k$ and $m=|H|$. Consider the sequences
\begin{align*}
{\bf k}: &0, 1, -1,2, -2,\ldots, n/2,\\
{\bf \bar k}: & 0, 1, -2, 3, -4, \ldots, n-1,
\end{align*}
in $\mathbb{Z}_{2^k}$, where ${\bf \bar k}$ is a symmetric sequencing. By Theorem \ref{newseqgen}, there exists a sequence ${\bf g}: 0, g_1,\ldots, g_{mn-1}$ in $G$ such that both ${\bf g}$ and ${\bf \bar g}$ are permutations of $G$. Moreover, by \eqref{g11}, we have $g_1=(1,0) \in \mathbb{Z}_{2^k} \times H$ and $g_{mn-1}=(n/2,0)$, the unique element of order 2 in $G$. 

Let $\alpha=1+mn/2$. Since $\gcd(\alpha, mn)=1$, the map $g \mapsto \alpha g$ is an isomorphism of $G$. Consider the sequence:
$$g_1,g_2,\ldots, g_{mn-1}, \alpha g_1,\alpha g_2, \ldots, \alpha g_{mn-1}.$$
Each element of $G \backslash \{0_G\}$ appears twice in the sequence. Moreover, the consecutive differences are:
$$-\alpha g_{mn-1}+g_1,\bar g_2, \bar g_3, \ldots, \bar g_{mn-1}, -g_{mn-1}+\alpha g_1,\alpha \bar g_2, \ldots, \alpha \bar g_{mn-1}.$$
To show that this sequence also contains every element of $G\backslash \{0_G\}$ twice, it is sufficient to show that $-g_{mn-1}+\alpha g_1=g_1$ and $-\alpha g_{mn-1}+g_1=\alpha g_1$. Both of these equations follow from $(\alpha-1)g_1=(mn/2) g_1=(mn/2)(1,0)=(n/2,0)=g_{mn-1}$ in $G$ (since $m$ is odd, $mn/2=n/2 \pmod{n}$).
\end{proof}

%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Anderson's and Bailey's conjectures}\label{3}

M\"{u}esser and Pokrovskiy \cite{MP} used techniques from probabilistic combinatorics to prove that all sufficiently large groups are sequenceable. To implement their results, we have the following definition based on their notation. 

\begin{definition}
Let $G$ be a group and $A,B\subseteq G$ such that $|A|=m$ and $|B|=m-1$. A {\it rainbow Hamilton sequence} in $K^+[A;B]$ is a sequence $g_1,g_2,\ldots, g_m$ such that $A=\{g_1,\ldots, g_{m}\}$, and $B=\{g_1g_2,\ldots, g_{m-1}g_m\}$. A {\it rainbow Hamilton sequence} in $K^-[A;B]$ is a sequence $g_1,g_2,\ldots, g_m$ such that $A=\{g_1,\ldots, g_{m}\}$, and $B=\{g_1^{-1}g_2,\ldots, g_{m-1}^{-1}g_m\}$. In both cases, the sequence is said to be from $g_1$ to $g_{m}$. 
\end{definition} 

Let $\pi: G \rightarrow G^{ab}$ be the abelianization projection, where $G^{ab}=G/[G,G]$ and $[G,G]$ is the commutator subgroup. For a subset $X=\{x_1,\ldots, x_k\} \subseteq G$, we write $\sum X$ to denote the projection $\pi(x_1x_2,\ldots x_k) \in G^{ab}$ (which is independent of the order of the $x_i$'s). For convenience, instead of writing $\pi(a)+\pi(b)$ or $\pi(a)-\pi(b)$, we simply write $a+b$ and $a-b$. 

The results in this section are basic corollaries of the following theorem, which is restated here for convenience. 

\begin{theorem}\label{mpt}
{\rm \cite[Thm.\ 6.9]{MP}} Let $G$ be a sufficiently large group of order $n$. Let $V,C \subseteq G$ and $x,y \in G$ such that $|V|+1=|C| \geq n-\sqrt{n}$, $x\neq y$, and further suppose that $1_G \notin C$ if $G$ is an elementary abelian $2$-group. 

\begin{itemize}
\item[i)] If $\sum C=y-x$, then there exists a rainbow Hamilton sequence from $x$ to $y$ in $K^-[\{x,y\} \cup V;C]$.

\item[ii)] If $\sum C=x+y+2 \sum V$, then there exists a rainbow Hamilton sequence from $x$ to $y$ in $K^+[\{x,y\} \cup V;C]$.
\end{itemize}
\end{theorem}

The following corollary supports Bailey's conjecture \cite{B} that all finite groups, except the elementary 2-groups of order at least 4, are 2-sequenceable.

 \begin{cor}\label{2s}
All sufficiently large groups, except elementary $2$-groups, are $2$-sequenceable.
\end{cor}

\begin{proof}
Since $G$ is not an elementary 2-group, there exists $g\in G$ such that $g^2 \neq 1_G$. Let $x=g$, $y=g^2$, $V=G \setminus \{1_G,g,g^2\}$, $C=G \setminus \{1_G,g^{-1}\}$. One has
$$\sum C=\sum G-(-g)=g=y-x.$$
By Theorem \ref{mpt}, there exists a rainbow Hamilton sequence $g_1,g_2,\ldots, g_k$ from $g_1=x=g$ to $g_k=y=g^2$ in $K^-[\{x,y\} \cup V;C]$ which yields the 2-sequencing $1_G=g_0,g_1,g_2,\ldots, g_k$ in $G$. To see this, we note that $g_0,g_1,\ldots, g_k$ is a permutation of $G$, and moreover, each element of $G\setminus \{g^{-1}\}$ appears exactly once in the sequence of consecutive quotients $g_0,g_0^{-1}g_1,\ldots, g_{k-1}^{-1}g_{k}$ except for $g$ which appears twice (while $g^{-1}$ does not appear).\end{proof}

The following corollary supports Anderson's conjecture \cite{Anderson} that all binary groups, except the quaternion group, are symmetric sequenceable.

\begin{cor}\label{ss}
All sufficiently large binary groups are symmetric sequenceable. 
\end{cor}

\begin{proof}
Let $H$ be a Sylow 2-subgroup of $G$. Then $H$ is either isomorphic to the cyclic group $\mathbb{Z}_{2^m}$, $m\geq 1$, or isomorphic to the dicyclic group $Q_{2^m}$, $m\geq 3$. 

A binary group $G$ is symmetric sequenceable if and only if $G/\mathbb{Z}_2$ is 2-sequenceable \cite{Anderson3}, where $\mathbb{Z}_2$ is the normal subgroup of $G$ generated by the unique element of order 2. There are two cases:

Case 1. $H$ is isomorphic to $\mathbb{Z}_{2^m}$ for some $m\geq 1$. In this case, the Sylow 2-subgroup of $G/\mathbb{Z}_2$ is isomorphic to $\mathbb{Z}_{2^{m-1}}$. Since $G/\mathbb{Z}_2$ is abelian and its Sylow 2-subgroup is nontrivial and cyclic ($m>1$ for $G$ large enough), then $G/\mathbb{Z}_2$ is sequenceable, hence 2-sequenceable. 

Case 2. $H$ is isomorphic to $Q_{2^m}$, $m\geq 3$. In this case $G/\mathbb{Z}_2$ is a dihedral group and non-abelian for $m$ large enough, hence sequenceable. 

In both cases $G/\mathbb{Z}_2$ is 2-sequenceable. It follows that $G$ is symmetric sequenceable.
\end{proof}

Next, we discuss supersequenceable groups. Let $G$ be a finite group and $[G,G]$ be the commutator subgroup of $G$. Let $g_1,\ldots, g_n$ be an ordering of $G$ and $p=g_1g_2\cdots g_n$ and consider the coset $P(G)=p[G,G]$, which is independent of the choice of the original ordering. 

\begin{definition}A finite group $G$ is {\it supersequenceable} if for every $h\in p[G,G]$ there exists a sequence $g_1,\ldots, g_m$ such that $g=g_m$ and either ${\bf \bar g}$ is a sequencing or ${\bf \check g}$ is an R-sequencing in $G$. 
\end{definition} 

Supersequenceable groups include all abelian groups \cite{AKP,K2}, dihedral groups of order $4n+2\geq 10$ \cite{K2}, and dihedral groups of order $4q$, where $q$ is an odd prime. 

\begin{cor}\label{superseq}
All sufficiently large groups are supersequenceable. 
\end{cor}

\begin{proof}
Let $h\in P(G)$. Apply Theorem \ref{mpt} with $x=1_G$, $y=h$, $V=G \setminus \{1_G,h\}$, and $C=G\setminus \{1_G\}$. One verifies that
$$\sum C=\sum G=h=y-x,$$
which implies that there exists a rainbow Hamilton sequence $g_1,g_2,\ldots, g_k$ form $g_1=1_G$ to $g_k=h$ and the consecutive quotients $g_1, g_1^{-1}g_2,\ldots, g_{k-1}^{-1}g_k$ form a sequencing in $G$. Thus, $G$ is supersequenceable. 
\end{proof}
%%%%%%%%%%%%%%%%%
\section{Partial sequencings and double sequencings}\label{4}
Let ${\bf w}: w_1,\ldots, w_k$ be a sequence of distinct elements in $W \subseteq G$. We say ${\bf w}$ is a partial harmonious sequence in $W$ if the (linear) consecutive products $w_1w_2,\ldots, w_{k-1}w_k$ are all distinct. If there exist a sequence $w_{k+1},\ldots, w_m$ in $W$ such that the extended sequence $w_1,\ldots, w_m$ is a harmonious sequence in $W$, then we say that ${\bf w}$ is extendable in $W$. 

Similarly, if $w_1,\ldots, w_k$ is a sequence of distinct elements in $W\setminus \{1_G\}$, we say ${\bf \bar w}$ is a partial R-sequencing in $W$ if the (linear) consecutive quotients $w_1^{-1}w_2,\ldots, w_{k-1}^{-1}w_k$ are all distinct elements of $W\setminus \{1_G\}$. Finally, if there exist $w_{k+1},\ldots, w_m$ in $W$ such that $w_m^{-1}w_1,w_1^{-1}w_2,\ldots, w_{m-1}^{-1}w_m$ is an R-sequencing in $W$, we say that ${\bf \bar w}$ is extendable in $W$. 

When is it possible to extend a partial harmonious sequence or a partial R-sequencing in a group $G$ to a full harmonious sequence in $G$? One expects that if the length of the partial sequence is small enough (relative to the size of the group) and if the group satisfies the Hall-Paige condition, then the sequence is extendable. We first give an example of a non-extendable partial sequence of length $n+2$ in a group of order $4n$, where $n$ is arbitrarily large. 

\begin{example}
Let $H$ be a harmonious group of even order $n$ and let $G=H \times \mathbb{Z}_4$. Any harmonious sequence $g_1,\ldots, g_{4n}$ in $G$ is partitioned into blocks $B_0,\ldots, B_{k-1}$, where within each block $B_i$, the second components equal $i$ modulo 2. Since only the transition from one block to the next gives rise to an adjacent sum with an odd second component and there are $2n$ elements in $G$ with an odd second component, it follows that $k=n$. Moreover, $\sum_{i~even} |B_i|=2n$ and so $|B_0|+n-1\leq 2n$ which yields $|B_0|\leq n+1$. Let $h_1,\ldots, h_n$ be a harmonious sequence in $H$. Then
$$(h_1,0),(h_2,0),\ldots, (h_n,0), (h_1,2),(h_2,0),$$
is a partial harmonious sequence of length $n+2$ in $G$ and it would fit in $B_0$ (since all of the second components are even), hence it cannot be extended to a harmonious sequence in $G$. 

The same example works for the R-sequencing case, where one uses an R-sequencing $h_1,\ldots, h_n$ for $H$.
\end{example}

The next lemma shows that partial harmonious sequences and partial R-sequencings shorter than $\sqrt{|G|}$ are extendable in a sufficiently large group $G$ satisfying the Hall-Paige condition. 

\begin{lemma}\label{ext}
Let $G$ be a sufficiently large group of order $n$ and $W \subseteq G$ such that $\sum W=0$. Furthermore, suppose that $1_G \notin W$ if $G$ is an elementary abelian $2$-group. Then every partial harmonious sequence (partial R-sequencing) of length $k\leq |W|-n+\sqrt{n}$ is extendable to a harmonious sequence (R-sequencing) in $W$. 
\end{lemma}

\begin{proof} 
Let $w_1,\ldots, w_k$ be a partial harmonious sequence in $W$. 
Let $x=w_k$, $y=w_1$, $V=W \setminus\{w_1,\ldots, w_k\}$, and $C=W \setminus \{w_1w_2,\ldots, w_{k-1}w_k\}$. One has $|V|+1=|C|=|W|-k+1 \geq n-\sqrt{n}$ and 
\begin{align*}
\sum C & = \sum W +w_1+w_k- 2\sum_{i=1}^{k}w_i\\
 & =\sum W+w_1+w_k-2 \left (\sum W -\sum V \right )\\ 
& =x+y+2\sum V,
\end{align*}
since $\sum W=0$. By Theorem \ref{mpt}, there exists a rainbow Hamilton sequence $w_k,w_{k+1}, \ldots, w_m,w_{m+1}=w_1$, from $x=w_k$ to $y=w_1$ in $K_G^+[\{x,y\} \cup V; C]$. Then, $w_1,\ldots, w_m$ is a harmonious sequence in $W$. The proof for partial sequencings is similar. 
\end{proof}

The following corollary extends the corresponding abelian results \cite{DJ} to large groups. Recall that in a group $G$ satisfying the Hall-Paige condition, $\sum G=0$ and in a binary group $G$ where $\imath_G$ is the unique element of order 2, $\sum G=i_G$. 
\begin{cor}\label{bun}
Let $G$ be a sufficiently large group.
\begin{itemize}
\item[i)] If $G$ satisfies the Hall-Paige condition, then $G\backslash \{1_G\}$ is harmonious. 
\item[ii)] If $G$ is a binary group with a unique involution $\imath_G$, then $G\backslash \{\imath_G\}$ and $G\backslash \{1_G,\imath_G\}$ are harmonious. 
\end{itemize}
\end{cor}

The next lemma gives necessary and sufficient conditions for a large enough subset of a sufficiently large group to be doubly harmonious and doubly R-sequenceable. 

\begin{lemma}\label{dhr}
Let $G$ be a sufficiently large group of order $n$ and $W\subseteq G$ with $|W| \geq  n - \sqrt{n/2}$. Furthermore, suppose that $G$ is not an elementary abelian $2$-group. Then $W$ is doubly harmonious (doubly R-sequenceable) if and only if $2\sum W=0$.\end{lemma}

\begin{proof}

First, consider the partial harmonious case. We claim that there exist $v_1,v_2 \in W$ (not necessarily distinct) such that $v_2v_1 \in W$. Let $x\in W$ be arbitrary and note that $|W\cap xW^{-1}|=|W|+|xW^{-1}|-|W \cup xW^{-1}|\geq 2|W|-|G|>0$. So there exists $v_1 \in W\cap xW^{-1}$ such that $v_1=xv_2^{-1}$ for some $v_2 \in W$ yielding $v_2v_1=x \in W$. 

Let $t=0$ if $|W|$ is even and $t=1$ if $|W|$ is odd. Let $H=G\times \mathbb{Z}_2$ and $V=W \times \mathbb{Z}_2 \setminus \{(v_1,0), (v_2,1)\}$ and $C=W \times \mathbb{Z}_2 \setminus \{(v_2v_1,1-t)\}$. Then
\begin{align*}
(v_1,0)+(v_2,1)+2\sum V & =(x,1)+2\left (2\sum W,t \right )-2(x,1)=-(x,1),\\
\sum C& =\left (2\sum W, t \right )-(x,1-t)=-(x,1),
\end{align*}
in $H^{ab}=G^{ab} \times \mathbb{Z}_2$. It follows from Theorem \ref{mpt} that there is a rainbow Hamilton path $g_1, g_2,\ldots, g_{2k}$ in $K^+[\{(v_1,0),(v_2,1)\}\cup V;C]$ from $g_1=(v_1,0)$ to $g_{2k}=(v_2,1)$. 
We claim that $\tau(g_1),\ldots, \tau(g_{2k})$ is a doubly harmonious sequence in $W$, where $\tau: H \rightarrow G$ is the projection onto the first component. The first components in $g_1,\ldots, g_{2k}$ contain each $g\in W$ exactly twice. Moreover, each element of $W\backslash \{v_2v_1\}$ has appeared as a first component twice among $g_{i-1}g_{i}$, $2\leq i \leq 2k$, while $v_2v_1$ has appeared exactly once. Additionally, $v_2v_1=\tau(g_{2k}g_1)$, which completes the proof that the sequence is doubly harmonious in $G$. 

Second, consider the partial R-sequencing case. Without loss of generality, suppose $1_G \notin W$. We claim that there exist distinct $u_1,u_2 \in W$ such that $u_2^{-1}u_1 \in W$. Let $y\in W$ be arbitrary and note that $|W\cap yW|\geq 2|W|-|G|>0$. So there exists $u_1 \in W$ such that $u_1=yu_2$ for some $u_2 \in W$ yielding $u_2^{-1}u_1=y \in W$. 

Let $V'=W \times \mathbb{Z}_2 \setminus \{(u_1,0), (u_2,1)\}$, and $C'=W \times \mathbb{Z}_2 \setminus \{(u_2^{-1}u_1,1-t)\}$. Then
$$\sum C'=\left (2\sum W,t \right)-(y,1-t)=-(y,1)=(u_2,1)-(u_1,0),$$
in $H^{ab}$. By Theorem \ref{mpt}, there exists a rainbow Hamilton path $g_1,g_2,\ldots, g_{2k}$ from $g_1=(u_1,0)$ to $g_{2k}=(u_2,1)$ in $K^-[\{(u_1,0),(u_2,1)\cup V';C']$. The first components of this sequence contain each element of $W$ exactly twice. Moreover, each element of $W\setminus \{u_2^{-1}u_1\}$ appears exactly twice as a first component among $g_{i-1}^{-1}g_i$, $2\leq i \leq 2k$, while $u_2^{-1}u_1$ appears exactly once. Additionally, $u_2^{-1}u_1=\tau(g_{2k}^{-1}g_1)$, which proves that we have obtained a double R-sequencing of $W$. 
\end{proof}

\begin{theorem}\label{bun2}
Let $G$ be a sufficiently large group. Then $G$ is doubly sequenceable, doubly R-sequenceable, doubly harmonious, and doubly R-harmonious.
\end{theorem}

\begin{proof}
The abelian cases are treated in \cite{DJ,Jav} and Theorem \ref{drs}. All sufficiently large non-abelian groups are sequenceable, hence doubly sequenceable. The rest of the claims follow from Lemma\ \ref{dhr}, since $2\sum G=0$ for every finite group $G$.  
\end{proof}

\end{document}